The critical magnetic fields of a super-conductor at temperatures 4K and 8K are 11 mA/m and 5.5 mA/m respectively. The transition

temperature is approximately?


1
Expert's answer
2021-10-13T10:42:07-0400
"\frac{T_c^2-16}{T_c^2-64}=\frac{11}{5.5}\\T_c=10.6\ K"